dumbkid101:

Evaluate the quotient of 1.05 x 10^3 and 1 x 10^3: 1.05 x 10^9 1.05 x 10^6 1.05 x 10^0 1.05 x 10^-6

dumbkid101:

@AP

dude:

Quotient = division \(\large \frac{1.05\times 10^3}{1\times 10^3}\) If an exponent is below the fraction you can convert it by putting it in the numerator with a negative exponent =\(1.05\times 10^3\cdot 1\times 10^{-3}\)